OpenStudy (anonymous):

Need help solving 1/3X + 1/4 = -5

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Awolflover1 @tanya123 @mathstudent55

hero (hero):

\[\dfrac{1}{3}x - \dfrac{1}{4} = -5\]

hero (hero):

@lesliee23 how far did you get with solving this one?

OpenStudy (anonymous):

don't we need a common denominator

OpenStudy (anonymous):

@Hero Typo? You did a - instead of +

hero (hero):

\[\dfrac{1}{3}x + \dfrac{1}{4} = -5\] @lesliee23 Do you know the common denominator?

OpenStudy (anonymous):

12

hero (hero):

Yes, basically one, approach for solving this is to multiply both sides by the common denominator, which, in this case is 12.

hero (hero):

\[(12)\dfrac{1}{3}x + (12)\dfrac{1}{4} = (12)-5\] @lesliee23, What do you think we should do next?

OpenStudy (anonymous):